Which compound represents an ethylenic hydro carbo

PLEASE EXPLAIN ALSO.
Answer #1

A REALLY basic way to explain ozonolysis will be this:

R2 R4 R2 R4

The double bond breaks upjust in half, and the resulting product is the double bond with an oxygen.

With this explanation we can see that in order to have only ONE main product we need to use a reactant that has a DOUBLE bond just in half of the chain and both parts of the chain have to be equals:

Only option possible with that condition: D

CH3CH2CH=CHCH2CH3 + O3\rightarrow CH3CH2CH=O + O=CHCH2CH3 They are practically mirror images of each other so they are the same exact formula.